OpenStudy (anonymous):

The probability of winning a particular game at the Game Place is . Find the odds against winning the game. what?

OpenStudy (anonymous):

u didnt complete the question...

OpenStudy (anonymous):

oh sorry 1/4

OpenStudy (anonymous):

let me rewrite it

OpenStudy (anonymous):

hmmm .ok

OpenStudy (anonymous):

he probability of winning a particular game at the Game Place is 1/4. Find the odds against winning the game.

OpenStudy (anonymous):

the*

OpenStudy (anonymous):

complete it. you didn't mention the probability of winning the game and okay, the answer will be 1-(the probability of win)

OpenStudy (anonymous):

3/4

OpenStudy (anonymous):

@AbhimanyuPudi is right.... probability of not winning = 1 - (probability of winning)

OpenStudy (anonymous):

okay so its 3/4?

OpenStudy (anonymous):

crap idk who to best respone

OpenStudy (anonymous):

@dianarock336 3/4 is the right answer.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

ok ..:)

OpenStudy (anonymous):

i know this is old and closed, but shouldn't this be, its 3 to 1

hartnn (hartnn):

yes, it must be 3/1 odds against = (1-p)/p p=1/4 here.